Use the order completeness property to show that the set S= {n/n+7:n belong to N} has a supremum and infimum
1
Expert's answer
2020-04-20T12:41:54-0400

The given set is "S=\{ \frac{n}{n+7} : n\in \N \}" .


i.e. "S=\{ \frac{1}{8},\frac{2}{9},\frac{3}{10},...........\}"

As "n<n+7 \ , \ \forall \ n\in \N"

"\implies \frac{n}{n+7} <1" ",\forall \ n\in \N" .

Therefore 1 is an upper bound of "S" .

Again by completeness property of "\R" ,every non empty subset of real number that has an upper bound also has a supremum in "\R" .

Claim : "\text{Supremum of S i,e } \text{Sup}\ S=1" .

We know that An upper bound "u" of a non empty set "S" in "\R" is the Supremum of "S" iff for every "\epsilon>0" there is an "x\in S" such that

"u-\epsilon <x" .

As "\frac{n}{n+7}=1-\frac{7}{n+7}" .So for every "\epsilon >0" we find a "n\in N" such that


"\frac{7}{n+7}<\epsilon" .

Hence , for every "\epsilon>0" there is an element of "x\in S" such that

"1-\epsilon <1-\frac{7}{n+7} =\frac{n}{n+7}=x."

Hence 1 is an least upper bound of S.

As "8n\geq n+7 , \forall \ n\in \N \implies \frac{n}{n+7}\geq \frac{1}{8} , \forall \ n\in \N" .

Therefore "\frac{1}{8}" is a Lower bound of S.

Therefore , by Completeness property of "\R" , "S" has infimum .

As "\frac{1}{8}" is an element of S ,so by definition of infimum "\text{inf} \ S =\frac{1}{8}."